# Parameterized Prompt for Incremental Object Detection

### v2

@@ Abstract
Recent studies have demonstrated that incorporating trainable prompts into pretrained models enables effective incremental learning. However, the application of prompts in incremental object detection (IOD) remains underexplored. Existing prompts pool based approaches assume disjoint class sets across incremental tasks, which are unsuitable for IOD as they overlook the inherent co-occurrence phenomenon in detection images. In co-occurring scenarios, unlabeled objects from previous tasks may appear in current task images, leading to confusion in prompts pool. In this paper, we hold that prompt structures should exhibit adaptive consolidation properties across tasks, with constrained updates to prevent catastrophic forgetting. Motivated by this, we introduce Parameterized Prompts for Incremental Object Detection (P2IOD). Leveraging neural networks global evolution properties, P2IOD employs networks as the parameterized prompts to adaptively consolidate knowledge across tasks. To constrain prompts structure updates, P2IOD further engages a parameterized prompts fusion strategy. Extensive experiments on PASCAL VOC2007 and MS COCO datasets demonstrate that P2IOD’s effectiveness in IOD and achieves the state-of-the-art performance among existing baselines.
@@ 1 Introduction
In response to external changes, humans possess strong adaptability, allowing them to incrementally accumulate knowledge. Similarly, we expect object detection algorithms to learn in an incremental manner, a task termed incremental object detection (IOD). However, existing detection methods suffer from catastrophic forgetting [18] during incremental learning. This issue arises because current detection frameworks rely on predefined labeled datasets [37], implicitly assuming static data distributions. When learning from dynamically data distributions, these frameworks tend to forget previously learned knowledge [26], resulting in severe performance degradation on previous tasks.
@@
To address this challenge, many methods [7, 1] leverage the inherent co-occurrence phenomenon in IOD, where detection images typically contain both labeled objects from the current task and unlabeled objects from previous tasks. In such co-occurring scenarios, object distribution remains relatively static [2], providing latent knowledge to supplement previous tasks. Therefore, a key issue in IOD is properly leveraging the static distribution of objects in co-occurring scenarios. Recently, with the rise of pre-trained models, prompting has emerged as a promising direction for incremental learning. However, whether prompting is suitable for IOD’s co-occurring scenarios remains unexplored. Gaurav et al. [3] first introduce prompting into IOD, adopting a well-established prompts pool from incremental classification. We observe that the prompts pool exhibits confusion when incorporating the knowledge of the static object distribution in co-occurring scenarios, leading to a negative impact on performance.
Figure 1: Heatmap showing similarity weights between objects and task-specific prompts after four incremental learning steps. Gray indicates irrelevance, red indicates positive correlation, and blue indicates negative correlation. Due to the co-occurrence phenomenon, task 1 objects exhibit high similarity not only with their corresponding task prompt but also with prompts from other tasks.
@@
An ideal prompts pool stores task-specific prompts learned from different tasks and matches objects to its most relevant prompts based on similarity weight during inference [29]. However, when leveraging the static distribution of objects in co-occurring scenarios, the prompts pool encounters severe confusion, specifically manifesting as matching and task confusion. The former matching confusion refers to an object that cannot match the most relevant prompt. As shown in Fig. 1, we visualize the similarity weight between an object and different prompts. It can be observed that since the object appear across all tasks, they exhibit high similarity with all task-specific prompts, making it impractical to match the most relevant prompt. On the other hand, the task confusion refers to task-specific prompts learning knowledge outside of its tasks. The unlabeled previous objects in co-occurring scenarios provide latent knowledge, causing the prompts learned for the current task to incorporate knowledge from all previous tasks, which undermines the clarity of the prompt’s representation. We refer to the matching and task confusion introduced by the prompts pool in IOD as prompts pool confusion, which negatively affects IOD’s performance.
@@
To tackle the above problems, this paper proposes Parameterized Prompt for Incremental Object Detection (P2IOD). We argue that preserving prompts in a task-isolated manner leads to confusion when handling co-occurring scenarios in IOD. To address this limitation, we propose a novel framework for prompt-based IOD. In our framework, the structure for preserving prompt knowledge exhibits an adaptive consolidation property, ensuring that knowledge is preserved holistically across tasks while dynamically updating previously learned knowledge according to the co-occurring objects in the current task. Moreover, the prompt structure constrains updates to critical parameters, mitigating catastrophic forgetting. Building upon this idea, P2IOD redesigns the prompts pool as a parameterized multi-layer perceptron (MLP) to generate prompts, so as to exploit the adaptive consolidation property of neural networks, which naturally update learned knowledge in response to losses from co-occurring objects. We further interpret the constraint on parameterized prompts as a form of model fusion, where the parameters of previous and current prompts are preserved or merged based on their importance and consistency, ensuring that the knowledge from each task is retained. In addition, we introduce pseudo-labeling during training to mine latent knowledge from co-occurring objects.
@@
(i) This is the first work to investigate the issue of prompts pool confusion caused by the co-occurrence phenomenon to the best of our knowledge.
(ii) We proposed a novel framework for prompt-based IOD, emphasizing that prompt structures should exhibit adaptive consolidation properties across tasks, with constrained updates to prevent catastrophic forgetting.
(iii) We propose P2IOD, which redesigns the prompts pool as parameterized prompts and employs parameterized prompt fusion to constrain parameter updates.
(iv) Extensive experiments on PASCAL VOC2007 and MS COCO datasets demonstrate the effectiveness of the proposed method in IOD, achieving state-of-the-art performance in existing baselines.
Figure 2: The overall framework of P2IOD. To address the issue of prompt pool confusion, P2IOD redesigns the prompt pool as a parameterized prompt structure consisting of multi-layer perceptron (MLP) bottlenecks. P2IOD introduces independent parameterized prompts at each decoder layer to ensure the diversity of prompts. To further alleviate the problem of catastrophic forgetting, P2IOD proposes a parameterized prompt fusion mechanism, which adds an additional fusion process after each incremental training process to better preserve task information.
@@ 2.2 Incremental Object Detection
The distinction between incremental object detection and other incremental tasks lies in the co-occurrence phenomenon inherent in detection scenarios. Co-occurring scenarios contain numerous unlabeled objects from previous tasks that can supplement previous task knowledge. Knowledge distillation [10] provides a flexible way to mine previous task knowledge. Such approaches [23, 7, 4] employ the original detector to regularize the outputs and intermediate features of the incremental detector, thereby facilitating the transfer of knowledge in training data from the original to the incremental detector. As this knowledge inherently contains information about unlabeled objects, it enables the implicit mining of unlabeled objects within the co-occurring scenarios. Furthermore, some methods [13, 3] explicitly mine unlabeled objects through pseudo-labeling. These methods use the original detector to label the objects in the training data and subsequently filter out incorrect labels based on specific criteria. The static distribution labels obtained through the pseudo-labeling method allow the detectors to be immune to catastrophic forgetting. The above methods exhibit strong performance in co-occurring scenarios.
Recently, with the rise of PTM, the prompting of the PTM has become a promising direction for incremental learning. Prompting of the PTM stores prompts as an additional memory module, allowing the PTM to learn and retain relevant information from each incremental task. Gaurav [3] constructs a prompts pool to store the prompts learned from different tasks and matches the most relevant prompt during inference. However, we discover that introducing a prompts pool faces severe prompts pool confusion in co-occurring scenarios. Therefore, effectively incorporating prompts into PTM still requires further research in IOD.
@@ 3.1 Object Detection Baseline
We introduce parameterized prompts on the transformer-based Deformable-DETR [34] and Co-DETR [36] to validate our research motivation. In transformer-based object detection frameworks, there exist two types of attention mechanisms. The first is the multi-head attention mechanism [25] in Transformers. Given a query element and a set of key elements, the multi-head attention module obtains attention weights based on the similarity between the query-key pairs and adaptively aggregates important features according to the attention weights. To enable the model to focus on content from different representational subspaces and different positions, the multi-head attention mechanism combines the outputs of multiple attention heads with different learnable weights. Let $q\in{\Omega_{q}}$ indexes a query element with representation feature ${z_{q}}\in{\mathbb{R}^{C}}$, and $k\in{\Omega_{k}}$ indexes a key element with representation feature ${x_{k}}\in{\mathbb{R}^{C}}$, where $C$ is the feature dimension, $\Omega_{q}$} and $\Omega_{k}$} specify the query and key elements, respectively. The attention weights $A_{mqk}$} are calculated by
@@
The second attention mechanism is the deformable attention mechanism [34]. This design not only preserves the spatial structure of the feature map but also helps the detector accelerate convergence and reduce computational complexity. Given an input feature map $x\in{\mathbb{R}^{C\times H\times W}}$, let $q$ index a query element with content feature $z_{q}$} and a 2-d reference point $p_{q}$, the deformable attention feature is calculated by
@@
The multi-head attention mechanism’s global attention ensures that features interact fully with the added prompts. In contrast, in the deformable attention mechanism, each feature interacts only with a limited number of positions, making it difficult to focus on the information in the additional prompts. Therefore, we introduce the prompts only in the multi-head attention mechanism, which is used for object query interaction in the decoder.
@@ 4.1 Overview
We propose P2IOD to avoid the confusion in prompt-pool-based IOD methods when learning co-occurring object knowledge. We hold that prompt structures should exhibit the ability to adaptively consolidate knowledge across tasks while constraining updates to prevent catastrophic forgetting. Therefore, P2IOD redesigns the prompts pool into parameterized prompts, leveraging neural networks’ inherent adaptive consolidation to learn new knowledge while selectively updating previous knowledge from co-occurring scenarios. The parameterized prompts are implemented as multi-layer perceptron (MLP) bottlenecks composed of feedforward networks and are integrated into different decoder layers to enhance prompt diversity. To constrain updates to the prompt structure, P2IOD proposes a parameterized prompt fusion strategy. Moreover, similar to the approach in [3], P2IOD also incorporates pseudo-labeling to mine latent knowledge of co-occurring objects in the scene. During incremental training, only the parameters of class embeddings, bounding box embeddings, and the parametrized prompt structure ($\theta$) are trainable, while all other parameters ($\theta^{*}$) remain frozen to prevent knowledge forgetting. Fig. 2 illustrates the complete framework.
@@ 4.2 Parameterized Prompt Structure
We hold that the prompt structure should adaptively consolidate the latent knowledge that emerges in the co-occurring scenarios. To achieve this, we design the prompt structure as an MLP bottleneck composed of FNN layers rather than the prompts pool. This design encodes prompt-related knowledge into the neural network weight space and generates instance-specific prompts.
We follow the method in [3] by employing the frozen pre-trained detector as a query function to extract queries, which are then utilized as inputs to the MLP bottleneck. Given an input instance $x$, a set of proposals $P\in{\mathbb{R}^{N\times D}}$ is generated through a single pass of $P={\theta^{*}}\left(x\right)$, where $N$ is the number of proposals and $D$ is the embedding dimension of each proposal. $P$ contains the instance-related information preliminarily extracted by the frozen pre-trained detector for the instance $x$. However, the number of $P$ is too large to be directly used as queries. Following [3], we introduce a simple averaging operation to compress all proposals, resulting in $P\in{\mathbb{R}^{1\times D}}$. The entire query function $Q$ can be represented as follows:
@@
${\text{MH}}{{\text{A}}_{p}}\left({{q_{o}},p}\right)=\sum\limits_{m=1}^{M}{{W_{m}}\left[{\sum\limits_{k}{{A_{mqk}}\cdot{{{W^{\prime}}_{m}}\left[{q_{o}}:{p_{v}}\right]}}}\right]},$ (6)
${A_{mqk}}={\text{softmax}}\left({\frac{{q_{o}^{T}U_{m}^{T}{{V_{m}}\left[{q_{o}}:{p_{k}}\right]}}}{{\sqrt{{C_{v}}}}}}\right),$ (7)
where $q_{o}$}} is the objects queries in [34]; $\left[{x:y}\right]$ represents the concatenate operation. Following [28], we assign $p\in{\mathbb{R}^{{L_{p}}\times D}}$ into ${p_{k}}\in{\mathbb{R}^{\frac{{{L_{p}}}}{2}\times D}}$ and ${p_{v}}\in{\mathbb{R}^{\frac{{{L_{p}}}}{2}\times D}}$, and concatenate them to ${V_{m}}{q_{o}}$ and ${{W^{\prime}}_{m}}{q_{o}}$ respectively, while keeping $q_{o}^{T}U_{m}^{T}$} unchanged. This manner ensures that the input and output sequence lengths remain the same before and after integrating prompts. To increase prompt diversity, we introduce independent parameterized prompts into each decoder layer of the frozen pre-trained detector.

### v4

@@ Abstract
Recent studies have demonstrated that incorporating trainable prompts into pretrained models enables effective incremental learning. However, the application of prompts in incremental object detection (IOD) remains underexplored. Our study reveals that existing prompt-pool-based approaches assume disjoint class sets across incremental tasks, which are unsuitable for IOD as they overlook the inherent co-occurrence phenomenon in detection. In co-occurring scenarios, unlabeled objects from previous tasks may appear in current task images, leading to confusion in prompts pool. In this paper, we hold that prompt structures should exhibit adaptive consolidation properties across tasks, with constrained updates to prevent confusion and catastrophic forgetting. Motivated by this, we introduce Parameterized Prompts for Incremental Object Detection (P2IOD). Leveraging neural networks global evolution properties, P2IOD employs networks as the parameterized prompts to adaptively consolidate knowledge across tasks. To constrain prompts structure updates, P2IOD further engages a parameterized prompts fusion strategy. Extensive experiments on PASCAL VOC2007 and MS COCO datasets demonstrate that P2IOD’s effectiveness in IOD and achieves the state-of-the-art performance among existing baselines. Code is available at https://github.com/EMLS-ICTCAS/P2IOD.
@@ 1 Introduction
In response to external changes, humans possess strong adaptability, allowing them to incrementally accumulate knowledge [11, 13]. Similarly, we expect object detection algorithms to learn in an incremental manner. Existing detection methods suffer from catastrophic forgetting [26] during incremental learning. This issue arises because current detection frameworks rely on predefined labeled datasets [50], implicitly assuming static data distributions. When learning from dynamic data distributions, these frameworks tend to forget previously learned knowledge [35], resulting in severe performance degradation.
Figure 1: Existing prompts-pool-based methods store task-specific prompts learned from different tasks and match objects to their most relevant prompts based on similarity weights during inference. However, preserving knowledge in a task-isolated manner causes confusion in IOD. In contrast, our method redesigns the prompts pool as parameterized prompts to holistically preserve and update knowledge across tasks, thus mitigating confusion.
@@
To address this challenge, many incremental object detection (IOD) methods [7, 24, 17, 27, 33] exploit the inherent co-occurrence phenomenon, where detection images typically contain both labeled objects from the current task and unlabeled objects from previous tasks. In such co-occurring scenarios, object distribution remains relatively static [1], providing latent knowledge to supplement previous tasks. A key problem in IOD thus lies in effectively leveraging the static object distribution present in co-occurring scenarios. Recently, with the rise of pre-trained models [4], prompting has emerged as a promising direction for incremental learning [12, 19]. Yet, its suitability for IOD’s co-occurring scenarios remains unexplored. Gaurav et al. [2] first introduce prompting into IOD, adopting a well-established prompts pool from incremental classification. We observe that the prompts pool exhibits confusion when incorporating the knowledge of the static object distribution in co-occurring scenarios, leading to a negative impact on performance.
@@
An ideal prompts pool stores task-specific prompts learned from different tasks and matches objects to its most relevant prompts based on similarity weight during inference [38]. However, when leveraging the static distribution of objects in co-occurring scenarios, the prompts pool encounters severe confusion, specifically manifesting as matching and task confusion. As shown in Fig. 1, the former matching confusion refers to an object that cannot match the most relevant prompt. We visualize the similarity weight between an object and different prompts. It can be observed that since the object appears across all tasks, they exhibit high similarity with all task-specific prompts, making it impractical to match the most relevant prompt. On the other hand, the task confusion refers to task-specific prompts learning knowledge outside of its tasks. The unlabeled previous objects in co-occurring scenarios provide latent knowledge, causing the prompts learned for the current task to incorporate knowledge from all previous tasks, which undermines the clarity of the prompt’s representation. We refer to the matching and task confusion introduced by the prompts pool in IOD as prompts pool confusion, which negatively affects IOD’s performance.
@@
To tackle the above problems, this paper proposes Parameterized Prompt for Incremental Object Detection (P2IOD). We argue that preserving knowledge in a task-isolated manner leads to confusion when handling co-occurring scenarios in IOD. To overcome the confusion, we advocate that the structure for preserving prompt knowledge should exhibit an adaptive consolidation property, ensuring that knowledge is preserved holistically across tasks, while previous task knowledge can be dynamically updated in co-occurring scenarios. Building upon this, we further mitigate catastrophic forgetting by constraining updates to critical parameters within the prompt structure. Based on this insight, P2IOD redesigns the prompts pool as a parameterized multi-layer perceptron (MLP), so as to leveraging the adaptive consolidation property inherent in neural networks, which naturally update learned knowledge in response to losses from co-occurring objects. We interpret the constraint on parameterized prompts as a form of model fusion, where the parameters of previous and current prompts are preserved or merged based on their importance and consistency, ensuring that the knowledge from each task is retained. In addition, we introduce pseudo-labeling to mine latent knowledge from co-occurring objects.
@@
(i) This is the first work to investigate the prompts pool confusion caused by the co-occurrence phenomenon. We further advocate that prompt structures should exhibit an adaptive consolidation property and adopt constrained updates to prevent confusion and forgetting.
(ii) We propose P2IOD, which redesigns the prompts pool as parameterized prompts and employs parameterized prompt fusion to constrain parameter updates.
(iii) Extensive experiments on PASCAL VOC2007 and MS COCO datasets demonstrate the effectiveness of the proposed method in IOD, achieving state-of-the-art performance in existing baselines.
Figure 2: The overall framework of P2IOD. To address the issue of prompts pool confusion, P2IOD redesigns the prompts pool as a parameterized prompt structure consisting of multi-layer perceptron (MLP) bottlenecks. To further alleviate catastrophic forgetting, P2IOD proposes a parameterized prompt fusion strategy, which adds an additional fusion process after each incremental training process.
@@ 2.2 Incremental Object Detection
The distinction between incremental object detection and other incremental tasks lies in the co-occurrence phenomenon inherent in detection scenarios. Co-occurring scenarios contain numerous unlabeled objects from previous tasks that can supplement previous task knowledge. Knowledge distillation [10, 14, 23, 44, 39] provides a flexible way to mine previous task knowledge. Such approaches [31, 7, 3, 42, 43] employ the original detector to regularize the outputs and intermediate features of the incremental detector, thereby facilitating the transfer of knowledge in training data from the original to the incremental detector. As this knowledge inherently contains information about unlabeled objects, it enables the implicit mining of unlabeled objects within the co-occurring scenarios. Furthermore, some methods [17, 2, 41] explicitly mine unlabeled objects through pseudo-labeling. These methods use the original detector to label the objects in the training data and subsequently filter out incorrect labels based on specific criteria. The static distribution labels obtained through the pseudo-labeling method allow the detectors to be immune to catastrophic forgetting. The above methods exhibit strong performance in co-occurring scenarios.
With the rise of PTM, the prompting of the PTM has become a promising direction for incremental learning. Prompting of the PTM stores prompts as an additional memory module, allowing the PTM to learn and retain relevant information from each incremental task. Gaurav [2] constructs a prompts pool to store the prompts learned from different tasks and matches the most relevant prompt during inference. However, we discover that introducing a prompts pool faces severe prompts pool confusion in co-occurring scenarios. Therefore, effectively incorporating prompts into PTM still requires further research in IOD.
@@ 3.1 Object Detection Baseline
We introduce parameterized prompts on the transformer-based Deformable-DETR [47] and Co-DETR [49] to validate our motivation. In transformer-based object detection frameworks, there exist two types of attention mechanisms. The first is the multi-head attention mechanism [34] in Transformers. Given a query element and a set of key elements, the multi-head attention module obtains attention weights based on the similarity between the query-key pairs and adaptively aggregates important features according to the attention weights. To enable the model to focus on content from different representational subspaces and different positions, the multi-head attention mechanism combines the outputs of multiple attention heads with different learnable weights. Let $q\in{\Omega_{q}}$ indexes a query element with representation feature ${z_{q}}\in{\mathbb{R}^{C}}$, and $k\in{\Omega_{k}}$ indexes a key element with representation feature ${x_{k}}\in{\mathbb{R}^{C}}$, where $C$ is the feature dimension, $\Omega_{q}$} and $\Omega_{k}$} specify the query and key elements, respectively. The attention weights $A_{mqk}$} are calculated by
@@
The second attention mechanism is the deformable attention mechanism [47]. It not only preserves the spatial structure of the feature map but also helps the detector accelerate convergence and reduce computational complexity. Given an input feature map $x\in{\mathbb{R}^{C\times H\times W}}$, let $q$ index a query element with content feature $z_{q}$} and a 2-d reference point $p_{q}$, the deformable attention feature is calculated by
@@
To the best of our knowledge, there is currently no research integrating prompts into deformable attention mechanisms. The difficulty lies in the fact that deformable attention is a spatially local attention structure, while prompt interaction requires a global attention structure, making their integration challenging. Therefore, we only introduce prompts in the multi-head attention mechanism, which is used for object query interaction in the decoder.
3.2 Pseudo Labeling for Mining Potential Objects
In IOD, unlabeled previous task objects may appear in the background of current task images. Properly mining these previous task objects can significantly reduce forgetting, while treating these objects as background can lead to more severe forgetting. To make a fair comparison, we introduce the same pseudo-labeling method as MD-DETR [2] to mine the knowledge of unlabeled objects in the background.
In $T_{t}$, we employ the detector trained on $T_{t-1}$ to infer on each training sample, obtaining predictions: ${{\hat{y}}_{i}}=\left\{{{{\hat{s}}_{i}},{{\hat{b}}_{i}}}\right\}$. Here, ${\hat{s}}_{i}$}} represents the score for the highest-scoring category, and ${\hat{b}}_{i}$}} provides the bounding box coordinates for this prediction. Pseudo-labeling mechanism [5, 9, 24] sets a threshold $\tau$ to filter predictions with ${\hat{s}}_{i}$}} higher than this threshold as pseudo label ${{\tilde{y}}_{i}}=\left\{{{{\tilde{c}}_{i}},{{\tilde{b}}_{i}}}\right\}$. The threshold $\tau$ ensures that only the most reliable predictions are used when generating pseudo labels. Here, ${\tilde{c}}_{i}$}} represents the pseudo label’s category name, ${\tilde{b}}_{i}$}} is the bounding box coordinates for this pseudo label. Pseudo labels incorporate the knowledge from previous tasks $\left\{{{T_{1}}\ldots{T_{t-1}}}\right\}$, effectively alleviating the detector’s forgetting.
@@ 4.1 Overview
We propose P2IOD to alleviate the confusion in prompt-pool-based IOD methods when learning co-occurring object knowledge. Fig. 2 illustrates the complete framework. We hold that prompt structures should exhibit the ability to adaptively consolidate knowledge across tasks while constraining updates to prevent catastrophic forgetting. Therefore, P2IOD redesigns the prompts pool into parameterized prompts, leveraging neural networks’ inherent adaptive consolidation to naturally update learned knowledge in response to losses from co-occurring objects. The parameterized prompts are implemented as multi-layer perceptron (MLP) bottlenecks composed of feedforward networks and are integrated into different decoder layers to enhance prompt diversity. Considering that IOD algorithms are often deployed on resource-constrained scenarios such as edge devices [22, 21], P2IOD proposes a parameterized prompt fusion strategy to constrain updates to the prompt structure with low computational overhead. During incremental training, only the parameters of class embeddings, bounding box embeddings, and the parametrized prompt structure ($\theta$) are trainable, while all other parameters ($\theta^{*}$) remain frozen to prevent knowledge forgetting. Fig. 2 illustrates the complete framework.
@@ 4.2 Parameterized Prompt Structure
We hold that the prompt structure should adaptively consolidate the potential knowledge that emerges in the co-occurring scenarios. To achieve this, we design the prompt structure as an MLP bottleneck composed of FNN layers rather than the prompts pool. This parameterized prompt structure encodes prompt-related knowledge into the neural network weight space and generates instance-specific prompts.
We follow the method in [2] by employing the frozen pre-trained detector as a query function to extract queries, which are then utilized as inputs to the parameterized prompt. Given an input instance $x$, a set of proposals $P\in{\mathbb{R}^{N\times D}}$ is generated through a single pass of $P={\theta^{*}}\left(x\right)$, where $N$ is the number of proposals and $D$ is the embedding dimension of each proposal. $P$ contains both object and background information related to the instance $x$, which is preliminarily extracted by the frozen pre-trained detector. However, the number of proposals in $P$ is too large to be directly used as query features, requiring compression. Unlike the approach in [2], where only object-related proposals are compressed, we find that jointly compressing both object and background proposals in P2IOD is more effective. We believe this is because including background knowledge in the prompts helps the detector better distinguish between foreground and background. In contrast, [2] shows that adding background query features degrades its performance. This degradation stems from the knowledge retention and matching mechanism of the prompts pool, which limits the acquisition of background knowledge, indicating that the prompts pool structure is not suitable for IOD. Detailed analysis and experimental comparisons can be found in Appendix B.4. The entire query function $Q$ can be represented as follows:
@@
${\text{MH}}{{\text{A}}_{p}}\left({{q_{o}},p}\right)=\sum\limits_{m=1}^{M}{{W_{m}}\left[{\sum\limits_{k}{{A_{mqk}}\cdot{\left[{{W^{\prime}}_{m}q_{o}}:{p_{v}}\right]}}}\right]},$ (6)
${A_{mqk}}={\text{softmax}}\left({\frac{{q_{o}^{T}U_{m}^{T}{\left[{V_{m}q_{o}}:{p_{k}}\right]}}}{{\sqrt{{C_{v}}}}}}\right),$ (7)
where $q_{o}$}} is the objects queries [47], $\left[{x:y}\right]$ represents the concatenate operation. Following [37], we assign $p\in{\mathbb{R}^{{L_{p}}\times D}}$ into ${p_{k}}\in{\mathbb{R}^{\frac{{{L_{p}}}}{2}\times D}}$ and ${p_{v}}\in{\mathbb{R}^{\frac{{{L_{p}}}}{2}\times D}}$, and concatenate them to ${{W^{\prime}}_{m}}{q_{o}}$ and ${V_{m}}{q_{o}}$ respectively, while keeping $q_{o}^{T}U_{m}^{T}$} unchanged. This manner ensures that the input and output sequence lengths remain the same before and after integrating prompts. To increase prompt diversity, we introduce independent parameterized prompts into each decoder layer of the frozen pre-trained detector.